OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (ivettef365):

substitute y for 125,000

OpenStudy (anonymous):

wene i do that i get the wrong answer

OpenStudy (ivettef365):

you have x - 62,500 = 0.25(125000) x - 62,500 = 31,250

OpenStudy (ivettef365):

can you solve for x

OpenStudy (ivettef365):

what did you get ??

OpenStudy (anonymous):

i dont know waht to do after x-62,500=31,250

OpenStudy (ivettef365):

you add 62,500 to both sides to then you have x - 62,500 = 31,250 +62,500 =+62,500 =============== x = 93,750 so that is the answer

OpenStudy (anonymous):

ok thanks

OpenStudy (ivettef365):

yw :)